OpenStudy (anonymous):

f(x)=3/x & g(x)=3x which statement is true

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@cnknd

OpenStudy (anonymous):

@hba

jimthompson5910 (jim_thompson5910):

f(g(x)) = 3/(3x) = x g(f(x)) = 3(3/x) = 9/x

jimthompson5910 (jim_thompson5910):

oh made a typo, sry f(g(x)) = 3/(3x) = 1/x